OpenStudy (anonymous):

Solve for x, given the equation squaroot of x - 5 + 7 = 11. x = 21, solution is extraneous x = 21, solution is not extraneous x = 81, solution is extraneous x = 81, solution is not extraneos

OpenStudy (tkhunny):

You must communicate better than that and you are making efforts to do so. Excellent! Subtract 7 from both sides and then stop to think about it.

OpenStudy (anonymous):

\[\sqrt{x-5}\] + 7=11

OpenStudy (anonymous):

like minus 7 from \[\sqrt{x-5}\]

OpenStudy (tkhunny):

Please don't use "minus" like that. The word "subtract" actually means what you want. \(\sqrt{x-5} + 7 = 11\) Subtract 7 from each side. 7 - 7 = ?? 11 - 7 = ??

OpenStudy (anonymous):

\[\sqrt{x-5}=4\]

OpenStudy (tkhunny):

Perfect. Now the thinking part. If that square root is going to work, what do we know about "x-5"?

OpenStudy (anonymous):

i do not know to be completely honest, but wouldnt you square both sides of \[\sqrt{x-5} =4\]

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

You can do that, but it is MUCH wiser to think about it first. Answer my question! If the square root is to work, then \(x-5 \ge 0\;or\;x \;\ge \;5\) Does that make any sense?